OpenStudy (anonymous):

Nancy has bought 360 tropical fish for a museum display. She bought 8 times as many goldfish as angelfish. How many of each type of fish did she buy? A. 320 angelfish, 40 goldfish B. 160 angelfish, 200 goldfish C. 40 angelfish, 320 gold fish D. 20angelfish, 160 goldfish

OpenStudy (anonymous):

need help ASAP!!!!!!

OpenStudy (texaschic101):

a + g = 360 8a = g now sub 8a in for g in the 1st equation a + 8a = 360 9a = 360 a = 360/9 a = 40 now sub 40 in for a in either of the equations to find g 8a = g 8(40) = g 320 = g She bought 40 angelfish and 320 goldfish

OpenStudy (texaschic101):

any questions ?

OpenStudy (anonymous):

i have another one

OpenStudy (anonymous):

Allan and Dave bowl together and their combined total score for one game was 375 points. Allan’s score was 60 less than twice Dave’s. What were their scores? Which is a system of equations to model the problem if x represents Dave’s score and y represents Allan’s score? A. x+y=60 y=2x-375 B. x + y = 375 y = 2x – 60 C. x + y = 375 y = 2x + 60 D. x – y = 375 y = 2x – 60

OpenStudy (texaschic101):

x + y = 375 y = 2x - 60 answer : B

OpenStudy (anonymous):

Milo wants to make a mixture that is 50% lemon juice and 50% lime juice. How much 100% lemon juice should he add to a juice that is 20% lemon juice and 80% lime juice to make 4 gallons of the 50% lemon/50% lime juice mixture? A. 0.5 gallon B. 2 gallons C. 1.5 gallons D. 2.5 gallons

OpenStudy (texaschic101):

I do not know how to do this one...sorry :(

OpenStudy (anonymous):

its ok

OpenStudy (texaschic101):

you got anymore besides the last one ?

OpenStudy (anonymous):

no

OpenStudy (texaschic101):

I hope I helped you even a little bit :)

OpenStudy (anonymous):

ok thx